OpenStudy (anonymous):

simplify the the following expression square root of -104 and the square root of -8?

OpenStudy (anonymous):

and as in plus? or two different numbers?

OpenStudy (anonymous):

there too differnet numbers

OpenStudy (anonymous):

needing two different answers as well I'd assume

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

\[\sqrt{-108}=\sqrt{4\times 26\times -1}=\sqrt{4}\times\sqrt{26}\times \sqrt{-1}=2\sqrt{26}i\]

OpenStudy (anonymous):

similarly \[\sqrt{-8}=2\sqrt{2}i\]

OpenStudy (anonymous):

-108 did u meane -104?